OpenStudy (anonymous):

Simplify the radical. radical 8+radical 50

OpenStudy (anonymous):

. \[\sqrt{8} = \sqrt{4} * \sqrt{2} -> 2(\sqrt{2}) \] \[\sqrt{50} = \sqrt{25} * \sqrt{2} -> 5(\sqrt{2})\] \[2(\sqrt{2}) + 5(\sqrt{2}) = 7(\sqrt{2}) \]